# Online Joint Fine-tuning of Multi-Agent Flows

### v1

@@ 1 Introduction
In this writeup I provide one approach for flow self-improvement based upon access to an environment simulator. [^1] This includes the typical machine learning setup of a dataset with gold labels, but is more general. This procedure might be invoked after hand-tuning of a specific flow yields diminishing returns; alternatively, more complex flows can be initially considered as the burden of hand-specification is reduced. The key idea is to reduce episode preferences to individual node output preferences by simulating the impact of changing the output of a single node. This requires access to an environment simulator, i.e., the ability to reset state to an intermediate point in a computation and evaluate alternative outputs. [^2] Thus the approach here is not single sample, although single sample approaches are highly desirable for self-improving during production deployment. Individual node output preferences can be optimized using parametric preference learning if all nodes are language models [4, 5, 6]; other non-language-model components can also utilize pairwise preference data, e.g., ranking systems [10].
@@
The reader should consider this document a white paper describing preliminary results, rather than a mature scientific publication. The primary purpose of this document is to facilitate specific replication of the Musique result. Although the approach is plausibly broadly applicable, systematic evaluation on a wider range of problems is necessary before drawing any strong conclusions.
Figure 1: Actual flow utilized for Musique-Answerable. Key realistic aspects are: dependent generation with loops (“retrieval” depends upon “query”, etc.); dynamic conversation state (“scratchpad”); conditional branching (“stop retrieval”); and lack of granular supervision.
@@ 2.1 The Challenges
Figure 1 on page 1 diagrams the actual flow utilized for the submission to the easier variant of the Musique leaderboard (“Musique Answerable”). A more detailed description is provided in Section 3. Here I highlight several realistic properties which create difficulties for flow learning. The first is looping, which makes it awkward to apply stagewise learning procedures: when a node is updated, this changes the distribution of inputs to that node. The second is dynamic conversation state and conditional branching, which cascades to different downstream inputs and outputs encountered by other nodes. Finally the lack of granular supervision implies only the result of the joint sequence of decisions of all nodes can be evaluated. These challenges are not unique to flow learning, but rather reflect the general difficulties of reinforcement learning.
@@ 3 Musique Results
  • Sufficiency (“Musique-Full” only): given the original query and the final scratchpad of documents, decide if the query could be answered given the database. The modification is diagrammed in Figure 3 on page 3.
@@
For Musique-Answerable, the Lora rank-8 Phi-based flow achieves a new state-of-the-art result at the time of this writing. As indicated in Table 1 on page 1, the untrained performance of this flow is non-zero, suggesting that the prompts initialize the flow near a highly rewarding region; but there are substantial benefits from fine-tuning.
Progressive Validation
Dataset Flow On Training Set Test Set
Answer F1 Support F1 Answer F1 Support F1
Musique-Ans Untrained 0.076 0.514 n/a n/a
Submitted Variant 0.729 0.904 0.693 0.92
Table 1: Fine-tuning improves performance. Although the untrained performance is quite poor, it is empirically sufficient to initialize the flow such that learning can succeed.
@@ 3.1 Hyperparameter Selection
The use of Coin Betting eliminates hyperparameters associated with optimization per se, but there are additional hyperparameters to be tuned. Musique is a large dataset and good performance is possible after a single pass over the training set [^4] Furthermore, in early experimentation, a second pass over the training set had no impact on performance on the validation set; therefore I never did more than one training pass., so I used progressive validation loss to make decisions when possible, as this allowed for early termination of experiments associated with bad ideas: see “Subsampling the Training Set” in 3.2 below.
Using periodic soft ref updating [6] with a larger DPO regularization parameter ($\beta=1$) was better than using a fixed reference policy and a lower DPO regularization parameter ($\beta=0.1$). I gated the soft ref update on progressive validation improvement in answer f1, because this seemed like a good idea; I have not yet tested whether unconditional soft ref updating yields better results.
I tried base models meta-llama/Meta-Llama-3-8B-Instruct [21], Qwen/Qwen1.5-32B-Chat [22], and microsoft/Phi-3-medium-128k-instruct [23]; the latter had the best performance. The best rank of the Lora adapter for different base models was determined by trying all values $\{2,4,8,16,32\}$; $8$ gave the best results for Phi. I did limited experimentation with models as large as meta-llama/Meta-Llama-3-70B-Instruct but such experiments ran slowly and did not exhibit substantial lift in progressive validation loss, so I focused on the smaller models. [^5] In my experience with fine-tuning flows for other problems, the performance advantage of large models can be substantial before training but is typically eliminated by fine-tuning.
@@ 3.2 Musique Specific Techniques
The training set of Musique contains a different distribution of “hops” (number of reasoning steps required to answer the question) than the dev and test sets, see Table 2 of [7]. Subsampling the training set to match these moments reduces training time by two-thirds and brings progressive validation loss on the training set more into correspondence with evaluation on the dev set. With subsampling, it takes an 8xH100 machine about a day to consume the answerable training set.
@@
With online learning, one can monitor progressive validation loss; consequently there is no need to have a distinct data split to prevent overfitting. Because data splits are inefficient, I chose to train on the dev set before submission. This yields improvement on the dev set (assessed progressively): this suggests both that additional training was not harmful and the dev set is self-redundant. Since I refused to abuse the submission oracle, I have no way to determine whether this additional training step was helpful or not, but I assumed it was and used it in the final submission. This assumption may be incorrect because the test sets are designed to be out-of-distribution with the training and dev sets.
Enforcing Grounded Answers
Musique is intended for grounded question-answering and not regurgitation of (potentially inaccurate) pre-trained knowledge. Therefore I postprocess the output of the flow to correspond to a span in one of the selected documents. This has no effect on support f1 or sufficiency metrics, but yields a slight improvement in answer f1 on the validation set (circa $0.01$ based upon limited testing).
@@
The maximum number of retrievals was capped at 4, because this was the minimum value which achieved near perfect support f1, and larger values ran more slowly.
4 Immediate Next Steps
In the near future I will update this document with results on the Musique-Full dataset, as well as some ablation studies. Furthermore, the flow generates a sequence of outputs of the Query Generation node as a nuisance parameter, which I’ll compare with the (unused for training) question decomposition metadata provided with the dataset.
@@ 5 Discussion
This work was inspired by several observations: 1) flows are SOTA architectures for multiple benchmark tasks of interest; 2) flows are currently mostly hand-engineered; 3) flows are typically short, i.e., the unrolled realized computation graph of a typical flow has 10 or less nodes in it; and 4) pre-trained and instruction-tuned foundation models induce reasonable prior distributions over actions using “obvious” prompts for initialization. This combination of factors motivates the application of model-free short-horizon online reinforcement learning techniques such as this one. If any of these factors are relaxed, we can expect the approach described here to exhibit multiple failure modes. As an extreme example, using a tabula rasa transformer model would presumably result in most rollouts yielding no episode preferences; and even if training were run long enough to generate successful rollouts, I suspect the procedure would ultimately overfit the training set and not generalize well. A more likely failure mode to be encountered in practice is a problem which is well-covered by pre-training (e.g., an automated software engineering flow) but requires a long execution chain to succeed (e.g., circa 50 node invocations); in this case, rollouts may still mostly fail to find improvements without a high-quality initialization of the flow; and the one-step-deviation credit assignment strategy (which corresponds to a local policy improvement guarantee) may not be powerful enough to converge to a strong result.

### v2

@@ 1 Introduction
In this writeup I provide one approach for flow self-improvement based upon access to an environment simulator, i.e., the ability to reset state to an intermediate point in the computation and evaluate alternative outputs. [^1] A simulator can typically be constructed from a machine-learning dataset, as in the Musique experiments. This procedure might be invoked after hand-tuning of a specific flow yields diminishing returns; alternatively, more complex flows can be initially considered as the burden of hand-specification is reduced. The key idea is to reduce episode preferences to individual node output preferences by simulating the impact of changing the output of a single node. Individual node output preferences can be optimized using parametric preference learning if all nodes are language models [4, 5, 6]; other non-language-model components can also utilize pairwise preference data, e.g., ranking systems [10].
@@
The reader should consider this document a white paper describing preliminary results, rather than a mature scientific publication. The primary purpose of this document is to facilitate replication of the state-of-the-art Musique result. Although the approach is plausibly broadly applicable, systematic evaluation on a wider range of problems is necessary before drawing any strong conclusions.
Figure 1: Actual flow utilized for Musique-Answerable. Key realistic aspects are: dependent generation with loops (the entire retrieval block is looped multiple times); dynamic conversation state (“scratchpad”); learned conditional branching (“stop retrieval”); and lack of granular supervision.
@@ 2.1 The Challenges
Figure 1 on page 1 diagrams the actual flow utilized for the submission to the easier variant of the Musique leaderboard (“Musique Answerable”). A more detailed description is provided in Section 3. Here I highlight several realistic properties which create difficulties for flow learning. The first is looping, which makes it awkward to apply stagewise learning procedures: when a node is updated, this changes the distribution of inputs to that node. The second is dynamic conversation state and conditional branching, which cascades to different downstream inputs and outputs encountered by other nodes. Finally the lack of granular supervision implies only the result of the joint sequence of decisions of all nodes can be evaluated. These challenges are not unique to flow learning and are idiosyncratic instantiations of the general difficulties of reinforcement learning.
@@ 3 Musique Results
  • Sufficiency (‘‘Musique-Full’’ only): given the original query and the final scratchpad of documents, decide if the query could be answered given the database. The modification is diagrammed in Figure 3 on page 3.
@@
For both Musique-Answerable and Musique-Full, the online fine-tuned flow achieves a new state-of-the-art result at the time of this writing. As indicated in Table 1 on page 1, the untrained performance of this flow is non-zero, suggesting that the prompts initialize the flow near a highly rewarding region; but there are substantial benefits from fine-tuning.
Progressive Validation Progressive Validation
Flow On Training Set On Dev Set Test Set
AnsF1 SuppF1 AnsF1 SuppF1 AnsF1 SuppF1
Untrained 0.085 0.492 0.074 0.437 - -
+Train on Train 0.729 0.904 0.629 0.887 - -
+Enforce Grounded - - 0.650 0.887 - -
+Train on Dev - - 0.761 0.92 0.693 0.92
Table 1: Assorted treatment effects for Musique Answerable. ‘‘--’’ indicates ‘‘not measured’’. Fine-tuning improves performance: Although the untrained performance is quite poor, it is empirically sufficient to initialize the flow such that learning can succeed. Enforcing grounded answers improves AnsF1: Post-processing the output of the flow to correspond to a span in a selected document is beneficial. Training on the Dev set is internally consistent: Progressive validation loss on the dev set improves, suggesting both additional training is not harmful and the dev set is self-redundant.
@@ 3.1 Hyperparameter Selection
The use of Coin Betting eliminates hyperparameters associated with optimization per se, but there are additional hyperparameters to be tuned. Musique is a large dataset and good performance is possible after a single pass over the training set [^3] Furthermore, in early experimentation, a second pass over the training set had no impact on performance on the validation set; therefore I never did more than one training pass., so I used progressive validation loss to make decisions when possible, as this allowed for early termination of experiments associated with bad ideas: see ‘‘Subsampling the Training Set’’ in 3.2 below.
Using periodic soft ref updating [6] with a larger DPO regularization parameter ($\beta=1$) was better than using a fixed reference policy and a lower DPO regularization parameter ($\beta=0.1$). I gated the soft ref update on progressive validation improvement in answer f1, because this seemed like a good idea; I have not tested whether unconditional soft ref updating yields better results.
I tried models mistralai/Mistral-7B-Instruct-v0.2 [21], meta-llama/Meta-Llama-3-8B-Instruct [22], Qwen/Qwen1.5-32B-Chat [23], and microsoft/Phi-3-medium-128k-instruct [24]; the latter had the best performance. The best rank of the Lora adapter for different models was determined by trying all values $\{2,4,8,16,32\}$; $8$ gave the best results for Phi. I did limited experimentation with models as large as meta-llama/Meta-Llama-3-70B-Instruct but such experiments ran slowly and did not exhibit substantial lift in progressive validation loss during the initial part of training, so I early terminated those experiments and focused on the smaller models. [^4] In my experience with fine-tuning flows for other problems, the performance advantage of large models can be substantial before training but is typically eliminated by fine-tuning.
@@ 3.2 Musique Specific Techniques
The training set of Musique contains a different distribution of ‘‘hops’’ (number of reasoning steps required to answer the question) than the dev and test sets, see Table 2 of [7]. Subsampling the training set to match these moments reduces training time by two-thirds and brings progressive validation loss on the training set more into correspondence with evaluation on the dev set. With subsampling, it takes an 8xH100 machine about a day to consume the answerable training set.
Enforcing Grounded Answers
Musique is intended for grounded question-answering and not regurgitation of (potentially inaccurate) pre-trained knowledge. Therefore I postprocess the output of the flow to correspond to a span in one of the selected documents. This has no effect on support f1 or sufficiency metrics, but yields an improvement in answer f1 on the validation set as indicated in Table 1 on page 1.
@@
With online learning, one can monitor progressive validation loss; consequently there is no need to have a distinct data split to prevent overfitting. Because data splits are inefficient, I chose to train on the dev set before submission. As indicated in Table 1 on page 1, this yields improvement on the dev set (assessed progressively): this suggests both that additional training was not harmful and the dev set is self-redundant. Without abusing the submission oracle, I have no way to determine whether this additional training step is helpful or not, but I assumed it was and used it in the final submission. This assumption may be incorrect because the test set is designed to be out-of-distribution with the training and dev sets.
Pairwise Inference
On the Full version of the dataset, the questions come in pairs, exactly one of which is answerable and the other is not. Enforcing this constraint (by sorting on the output of the sufficiency node in the flow) yields substantial improvement in sufficiency accuracy, which improves the combined f1+sufficiency leaderboard metrics. This is quantified in Table 2 on page 2.
@@
The maximum number of retrievals was capped at 4, because this was the minimum value which achieved near perfect support f1, and larger values ran more slowly. All base models were 4-bit double quantized using bitsandbytes [25, 26].
Progressive Validation Progressive Validation
Flow On Training Set On Dev Set Test Set
AnsF1 SuppF1 SuffAcc AnsF1 SuppF1 SuffAcc Ans+SuffF1 Supp+SuffF1
Untrained - - - 0.073 0.450 0.625 - -
+Train on Train 0.689 0.887 0.86 0.579 0.883 0.874 - -
+Enforce Grounded - - - 0.604 0.883 0.874 - -
+Pairwise Inference - - - 0.604 0.883 0.945 - -
+Train on Dev - - - 0.740 0.921 0.957 0.672 0.887
Table 2: Assorted treatment effects for Musique Full. ‘‘--’’ indicates ‘‘not measured’’. ‘‘Ans+SuffF1’’ (‘‘Supp+SuffF1’’) is computed by the leaderboard evaluation and combines AnsF1 (SuppF1, resp.) with Sufficiency Accuracy. Pairwise inference improves sufficiency accuracy: Enforcing the constraint that exactly one of each paired example is answerable is beneficial. Other treatment effects: See the caption of Table 1 on page 1 for explanation.
@@ 4 Discussion
This work was inspired by several observations: 1) flows are SOTA architectures for multiple benchmark tasks of interest; 2) flows are currently mostly hand-engineered; 3) flows are typically short, i.e., the unrolled realized computation graph of a typical flow has 10 or less nodes in it; and 4) pre-trained and instruction-tuned foundation models induce reasonable prior distributions over actions using ‘‘obvious’’ prompts for initialization. This combination of factors motivates the application of model-free short-horizon online reinforcement learning techniques such as this one. If any of these factors are relaxed, we can expect the approach described here to exhibit multiple failure modes. As an extreme example, using a tabula rasa transformer model would presumably result in most rollouts yielding no episode preferences; and even if training were run long enough to generate successful rollouts, I suspect the procedure would ultimately overfit the training set and not generalize well. A more likely failure mode to be encountered in practice is a problem which is well-covered by pre-training (e.g., an automated software engineering flow) but requires a long execution chain to succeed (e.g., circa 50 node invocations); in this case, rollouts may still mostly fail to find improvements without a high-quality initialization of the flow; and the one-step-deviation credit assignment strategy (which corresponds to a local policy improvement guarantee) may not be powerful enough to converge to a strong result.
The approach describe here requires simulator access, i.e., the ability to reset state to an intermediate point in the computation and evaluate alternative outputs. This is applicable both to offline tuning and some online production scenarios (e.g., a software development assistant could plausibly reset state and evaluate an alternative rollout); however, for self-improvement from production deployment, a single-sample approach is often required. For example, a website powered by a flow cannot ask the user to start over halfway through their session so that the algorithm can try something different and see if it works better. I’m currently investigating single-sample approaches.
There is a relative lack of submission activity for Musique-Full compared to Musique-Answerable. Why don’t researchers take their solutions for Musique-Answerable and also run them on Musique-Full? This highlights an important benefit of flow based architectures: if you want to add capabilities to an existing system over time, having specialized components is beneficial. Here, the difference between the two is simply adding a single node to the flow as in Figure 3 on page 3.